OpenStudy (cherie_magee):

Find the common difference for the sequence shown 1/4, 5/16, 3/8, ... 1/16 1/4 1/2

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

\[\frac{ 5 }{ 16}-\frac{ 1 }{ 4 }=\frac{ 5-4 }{ 16 }=\frac{ 1 }{ 16 }\]